OpenStudy (anonymous):

help with the equation 1/8*g +1=99

OpenStudy (anonymous):

(1/8)g=98 ==> g=89*8=712

OpenStudy (lalaly):

is it\[\frac{1}{8}g+1=99\]or\[\frac{1}{8g}+1=99\]

OpenStudy (anonymous):

It is the first one.

OpenStudy (lalaly):

then raheen got it :D

OpenStudy (lalaly):

wait wait

OpenStudy (lalaly):

\[\frac{1}{8}g=98\]\[g=98 \times 8\]\[g=784\]